Geometria e algebra lineare 2009/10

Alcune soluzioni per gli esercizi 10/2/2010

Esercizio 1. r = {(2, −t + 1, t) | t ∈ ℝ}. Esercizio 2. 1. {x + z = 2 | y = 1}

2. {(t, 0, t − 1) | t ∈ ℝ}. Esercizio 3. 1. {x_{1} + x_{2} + x_{3} = 1}.

2.

3. {(1 + t, 2 − t + u, 3 − u) | t, u ∈ ℝ}.